# Topologically protected negative entanglement

### v3

@@ I.1.2 Effect of bulk states on entanglement scaling due to topological EPs
$H(k,k_{y})=t\cos k_{y}\sigma_{x}+(a_{0}-t\sin k_{y})\sigma_{+}+((b_{0}-\cos k)^{B}+t\sin k_{y})\sigma_{-}\,,$ (S24)
@@
II.1 A. Exponential scaling of the topological gap with $L_{y}$
@@
$H_{\text{y-OBC}}(k)$ $=$ $\begin{pmatrix}h_{0}^{y}&h_{+}^{y}&\cdots&0&0\\ h_{-}^{y}&h_{0}^{y}&h_{+}^{y}&\cdots&0\\ 0&h_{-}^{y}&h_{0}^{y}&\cdots&0\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ 0&0&\cdots&h_{-}^{y}&h_{0}^{y}\end{pmatrix}_{L_{y}\times L_{y}}\text{with }h_{0}^{y}=\begin{pmatrix}0&a_{0}\\ (b_{0}-\cos k)^{B}&0\end{pmatrix},\,h_{+}^{y}=\begin{pmatrix}0&0\\ t&0\end{pmatrix},\,h_{-}^{y}=\begin{pmatrix}0&t\\ 0&0\end{pmatrix}.$ (S25)
@@
$\det(H_{\text{y-OBC}}(k))$ $=$ $\det(h_{0}^{y})\det\begin{pmatrix}h_{0}^{y}&h_{+}^{y}&\cdots&0\\ h_{-}^{y}&h_{0}^{y}&\cdots&0\\ \vdots&\ddots&\ddots&\vdots\\ 0&\cdots&h_{-}^{y}&h_{0}^{y}-h_{+}^{y}(h_{0}^{y})^{-1}h_{-}^{y}\end{pmatrix}=\det(h_{0}^{y})\det\begin{pmatrix}h_{0}^{y}&h_{+}^{y}&\cdots&0\\ h_{-}^{y}&h_{0}^{y}&\cdots&0\\ \vdots&\ddots&\ddots&\vdots\\ 0&\cdots&h_{-}^{y}&h_{0}^{y}\end{pmatrix}_{(L_{y}-1)\times(L_{y}-1)}$ (S26)
@@
$H^{\prime}(k)=Q^{-1}H_{\text{y-OBC}}(k)Q,\quad Q=\text{diag}\{1,r,r,r^{2},\cdots r^{L_{y}-1},r^{L_{y}-1},r^{L_{y}}\},\quad r=\sqrt{\left|\frac{(b_{0}-\cos k)^{B}}{a_{0}}\right|},$ (S27)
@@
$H^{\prime}(k)=\begin{pmatrix}0&te^{-ik_{y}}+t^{\prime}\\ te^{ik_{y}}+t^{\prime}&0\end{pmatrix},\quad\text{where }\quad t^{\prime}=\sqrt{a_{0}(b_{0}-\cos k)^{B}}.$ (S28)
Since a similarity matrix transform does not change the eigenspectrum, $H^{\prime}(k)$ possesses the same eigenvalues as $H_{\text{y-OBC}}$, with its bulk eigenvalues given by $E_{n}=\pm\sqrt{(te^{ik_{y}}+t^{\prime})(te^{-ik_{y}}+t^{\prime})}$. By substituting these values of $E_{n}$ into $\det[H_{\text{y-OBC}}(k)]=E_{e_{1}}(k)E_{e_{2}}(k)\prod_{n}E_{n}(k)$, we obtain
$\prod_{n}E_{n}(k)$ $=$ $-\prod_{k_{y}}(te^{ik_{y}}+t^{\prime})(te^{-ik_{y}}+t^{\prime})\quad\sim t^{2L_{y}}+c_{1}t^{2L_{y}-1}t^{\prime}+c_{2}t^{2L_{y}-2}t^{\prime 2}+\cdots+c_{L_{y}}t^{\prime 2L_{y}},$ (S29)
where in the RHS, the terms containing $e^{\pm ik_{y}}$ are incorporated into the coefficients $c_{1},c_{2},...,c_{L_{y}}$. As mentioned in the main text, the topological non-trivial condition is $|a_{0}(b_{0}-\cos k)^{B}|\leq t^{2}$, i.e. $|t^{\prime}|\leq|t|$. Consequently, the leading term in the RHS of the above equation is $t^{2L_{y}}$. Then using Eq. (S29), we have
$\det[H_{\text{y-OBC}}(k)]=E_{e_{1}}(k)E_{e_{2}}(k)*t^{2L_{y}}\left(1+O(t^{\prime}/t)\right)$ $=$ $\left(a_{0}(b_{0}-\cos k)^{B}\right)^{L_{y}}$
$E_{e_{1}}(k)E_{e_{2}}(k)$$\approx$ $\frac{\left(a_{0}(b_{0}-\cos k)^{B}\right)^{L_{y}}}{t^{2L_{y}}}.$ (S30)
@@
$\Delta=2|E_{e_{1}}|\sim\left(\frac{a_{0}(b_{0}-\cos k)^{B}}{t^{2}}\right)^{L_{y}/2}.$ (S31)
II.2 B. Scaling of entanglement entropy $S_{A}$ with system sizes $L$ and $L_{y}$ in the case of gapless flat bands
In this subsection, we derive the scaling relations of $S_{A}$ as a function of $L_{y}$ and $L$, specifically Eq. (14) in the main text:
$S_{A}\sim-\frac{1}{2}(BL_{y})^{2}\log L$ (S32)
@@
Below, we present a comprehensive derivation of Eq. (S33). Under OBCs in the $y$ direction, $k$ is still a good eigenstate label and the projector operator in $k$ space is defined as $P(k)=\sum_{n\in occ}\ket{\psi_{n}(k)^{R}}\bra{\psi_{n}(k)^{L}}$ |. Due to the NHSE which pushes all right eigenstates in the y-direction towards the same edge, they become highly edge-localized and hence almost orthogonal. This large overlap ensures that the corresponding left edge eigenstates exhibit large amplitudes and contribute most significantly to $P(k)$. Therefore, $P(k)$ is dominated by the edge state contributions:
$P(k)$ $=$ $\sum_{n\in occ}\ket{\psi_{n}(k)^{R}}\bra{\psi_{n}(k)^{L}}$ | (S34)
@@
In the following discussion, for the sake of brevity, we will omit the $k$ in |$\ket{\psi_{edge}(k)^{R}},\bra{\psi_{edge}(k)^{L}}$ |, rewriting them as |$\ket{\psi_{edge}^{R}},\bra{\psi_{edge}^{L}}$ |. To proceed, we note that $P_{edge}(k)$ is furthermore dominated by just one matrix element due to the exponential skin-localization of the edge states. To show this explicitly, we use the similarity transform [Eq. (S28)] to write the spatial profiles of the topological SSH edge states in the $y$ direction as
|$\ket{\psi^{\prime}_{edge}(y)}=\begin{pmatrix}\psi^{\prime A}(y)\\ \psi^{\prime B}(y)\end{pmatrix}\sim\begin{pmatrix}\epsilon^{y-1}\\ \epsilon^{L_{y}-y}\end{pmatrix},\quad\text{ where }\epsilon=\left|\frac{t^{\prime}}{t}\right|=\sqrt{\left|\frac{a_{0}}{t}(b_{0}-\cos k)^{B}\right|}.$ (S35)
@@
$|\psi_{edge}^{R}\rangle=Q|\psi^{\prime}_{edge}\rangle\sim\begin{pmatrix}\begin{pmatrix}1\\ r\epsilon^{L_{y}-1}\end{pmatrix}\\ r\begin{pmatrix}\epsilon\\ r\epsilon^{L_{y}-2}\end{pmatrix}\\ \vdots\\ r^{L_{y}-2}\begin{pmatrix}\epsilon^{L_{y}-2}\\ r\epsilon\end{pmatrix}\\ r^{L_{y}-1}\begin{pmatrix}\epsilon^{L_{y}-1}\\ r\end{pmatrix}\end{pmatrix},\quad\text{ and }\bra{\psi_{edge}^{L}}=Q^{-1}\bra{\psi^{\prime}_{edge}}\sim\begin{pmatrix}\begin{pmatrix}1\\ \epsilon^{L_{y}-1}/r\end{pmatrix}\\ \frac{1}{r}\begin{pmatrix}\epsilon\\ \epsilon^{L_{y}-2}/r\end{pmatrix}\\ \vdots\\ \frac{1}{r^{L_{y}-2}}\begin{pmatrix}\epsilon^{L_{y}-2}\\ \epsilon/r\end{pmatrix}\\ \frac{1}{r^{L_{y}-1}}\begin{pmatrix}\epsilon^{L_{y}-1}\\ 1/r\end{pmatrix}\end{pmatrix}^{T}.$ (S36)
@@
$P(k)\approx P_{edge}(k)=\ket{\psi_{edge}^{R}}\bra{\psi_{edge}^{L}}\sim\begin{pmatrix}\begin{pmatrix}1&\epsilon^{L_{y}-1}/r\\ r\epsilon^{L_{y}-1}&\epsilon^{2L_{y}-2}\end{pmatrix}&\frac{1}{r}\begin{pmatrix}\epsilon&\epsilon^{L_{y}-2}/r\\ r\epsilon^{L_{y}}&\epsilon^{2L_{y}-3}\end{pmatrix}&\cdots&\color[rgb]{0,0,1}\frac{1}{r^{L_{y}-1}}\begin{pmatrix}\epsilon^{L_{y}-1}&1/r\\ r\epsilon^{2L_{y}-2}&\epsilon^{L_{y}-1}\end{pmatrix}\\ r\begin{pmatrix}\epsilon&\epsilon^{L_{y}}/r\\ r\epsilon^{L_{y}-2}&\epsilon^{L_{y}-2}\end{pmatrix}&\begin{pmatrix}\epsilon^{2}&\epsilon^{L_{y}-1}/r\\ r\epsilon^{L_{y}-1}&\epsilon^{2L_{y}-4}\end{pmatrix}&\cdots&\frac{1}{r^{L_{y}-2}}\begin{pmatrix}\epsilon^{L_{y}}&\epsilon/r\\ r\epsilon^{2L_{y}-3}&\epsilon^{L_{y}-2}\end{pmatrix}\\ \vdots&\vdots&\vdots&\vdots\\ r^{L_{y}-1}\begin{pmatrix}\epsilon^{L_{y}-1}&\epsilon^{2L_{y}-2}/r\\ r&\epsilon^{L_{y}-1}\end{pmatrix}&r^{L_{y}-2}\begin{pmatrix}\epsilon^{L_{y}}&\epsilon^{2L_{y}-3}/r\\ r\epsilon&\epsilon^{L_{y}-2}\end{pmatrix}&\cdots&\begin{pmatrix}\epsilon^{2L_{y}-2}&\epsilon^{L_{y}-1}/r\\ r\epsilon^{L_{y}-1}&1\end{pmatrix}\end{pmatrix}_{2L_{y}\times 2L_{y}}$ (S37)
@@
The truncated projector $\bar{P}$ can be obtained by Fourier transforming each matrix element on the right-hand side (RHS) of Eq. (S37) i.e.
$\bar{P}=\begin{pmatrix}\bar{P}_{1,1}&\bar{P}_{1,2}&\cdots&\bar{P}_{1,2L_{y}}\\ \bar{P}_{2,1}&\bar{P}_{2,2}&\cdots&\bar{P}_{2,2L_{y}}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{P}_{2L_{y},1}&\bar{P}_{2L_{y},2}&\cdots&\bar{P}_{2L_{y},2L_{y}}\\ \end{pmatrix},\text{ where }\bar{P}_{m,n}\text{ is an {\hbox{l\times l}} matrix whose elements are }\bra{x_{1}}\bar{P}_{m,n}\ket{x_{2}}=\frac{1}{L}\sum_{k}e^{ik(x_{1}-x_{2})}P_{m,n}(k),$ (S38)
@@
with $b_{0}=1$ and the leading $k=\pm k_{1}=\pm\pi/L$ contributions substituted to obtain the 3rd line. The other matrix elements scale more slowly with $L$. Since $\bra{x_{1}}\bar{P}_{2L_{y},1}\ket{x_{2}}\sim O(1)$, we have established Eq. S33:
$\text{Tr}(\bar{P}^{2})\approx 2\text{Tr}(\bar{P}_{1,2L_{y}}\bar{P}_{2L_{y},1})\sim c_{1}L^{BL_{y}-1}+c_{2}L^{BL_{y}-1-2}+\cdots$ (S40)
@@
$S_{A}(p_{i})$ $=$ $-p_{i}\log p_{i}-(1-p_{i})\log(1-p_{i})=-cL^{\alpha}\log(cL^{\alpha})-(1-cL^{\alpha})\log(1-cL^{\alpha})$ (S41)
@@
$=$ $(cL^{\alpha}+1)\log(-1)-\log(cL^{\alpha})\approx-{\color[rgb]{0,0,1}\alpha\log L}+cL^{\alpha}\pi i-\log c.$
The other eigenvalue with $p_{i}^{\prime}=1-p_{i}$ (see Sect. III.B) contributes the same real part to $S_{A}$, but opposite imaginary part that cancels off, as numerically observed as twofold degenerate states. Therefore, the total entanglement entropy is dominated by the $\log L$ contribution in blue above which should be multiplied by 4, i.e.
$S_{A}=\sum_{p_{i}}-{p_{i}}\log p_{i}-(1-p_{i})\log(1-p_{i})\sim-4\times\left[\frac{BL_{y}-1}{2}+\left(\frac{BL_{y}-1}{2}-1\right)+\left(\frac{BL_{y}-1}{2}-2\right)+\cdots\right]\log L.$ (S42)
@@
$=$ $-\left\lceil\frac{B^{2}L_{y}^{2}-1}{2}\right\rceil\log L\,\sim-\frac{1}{2}(BL_{y})^{2}\log L,$
@@
Below, we further present the numerical verification of Eq. (S42). In Fig. S4(a), we take $L_{y}=3,B=1$, since $\frac{BL_{y}-1}{2}=1$, only the first order $p_{i}\sim L^{1}$ exists. The numerical fitting shown in Fig. S4(d) yields:
$S_{A}\sim-4.4072\log L\quad\approx-4\times\frac{BL_{y}-1}{2}\log L=-4\log L,$ (S44)
which demonstrates good agreement with Eq. (S42). In Fig. S4(b), where $L_{y}=5,B=1$, $\frac{BL_{y}-1}{2}=2$, both the first and second orders of $L$ in $p_{i}\sim L^{1}$, $p_{i}\sim L^{2}$ are present. We numerically obtain the scaling of entanglement entropy as depicted in Fig. S4(e):
$S_{A}\sim-11.7648\log L\approx-4\times\left(\frac{BL_{y}-1}{2}+\frac{BL_{y}-1}{2}-1\right)\log L=-12\log L.$ (S45)
For the case of $L_{y}=3,B=2$ shown in Fig. S4(c), the first, second and third orders of $L$ all contribute to the EE:
$S_{A}\sim-17.1980\log L\approx-4\times\left(\frac{BL_{y}-1}{2}+\frac{BL_{y}-1}{2}-1+\frac{BL_{y}-1}{2}-2\right)\log L=-18\log L.$ (S46)
@@
Figure S4: Excellent agreement of our super volume-law negative entanglement scaling result with numerics for various $B$ and $L_{y}$, both for the $\bar{P}$ eigenvalues $p_{i}$ [Eq. S33] (a-c) and the corresponding entanglement scaling $S_{A}\sim-\left\lceil(B^{2}L_{y}^{2}-1)/2\right\rceil\log L$ [Eq. S43]. Parameters: $t=0.5,a_{0}=2,b_{0}=1$.
II.3 C. Scaling of entanglement entropy $S_{A}$ for gapped flat bands
In the case of gapped bands where $b_{0}>1$ (e.g., $b_{0}=1.2$), there exists a small gap $\Delta$ (see Eq. (S31)) in the topological flat edge bands. Due to this gap, the entanglement entropy $S_{A}$ decreases with increasing system size $L$ and eventually stabilizes at a minimum constant value, $S_{min}$. In this section, we will demonstrate that:

### v4

@@ II.2 Detailed analysis of the topological flat band model
$H(k,k_{y})=t\cos k_{y}\sigma_{x}+(a_{0}-t\sin k_{y})\sigma_{+}+((b_{0}-\cos k)^{B}+t\sin k_{y})\sigma_{-}\,,$ (47)
@@
II.2.1 Exponential scaling of the topological gap with $L_{y}$.
@@
$H_{\text{y-OBC}}(k)$ $=$ $\begin{pmatrix}h_{0}^{y}&h_{+}^{y}&\cdots&0&0\\ h_{-}^{y}&h_{0}^{y}&h_{+}^{y}&\cdots&0\\ 0&h_{-}^{y}&h_{0}^{y}&\cdots&0\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ 0&0&\cdots&h_{-}^{y}&h_{0}^{y}\end{pmatrix}_{L_{y}\times L_{y}}$ (48)
with
$h_{0}^{y}=\begin{pmatrix}0&a_{0}\\ (b_{0}-\cos k)^{B}&0\end{pmatrix},\,h_{+}^{y}=\begin{pmatrix}0&0\\ t&0\end{pmatrix},\,h_{-}^{y}=\begin{pmatrix}0&t\\ 0&0\end{pmatrix}.$ (49)
@@
$\det(H_{\text{y-OBC}}(k))$ $=$ $\det(h_{0}^{y})\det\begin{pmatrix}h_{0}^{y}&h_{+}^{y}&\cdots&0\\ h_{-}^{y}&h_{0}^{y}&\cdots&0\\ \vdots&\ddots&\ddots&\vdots\\ 0&\cdots&h_{-}^{y}&h_{0}^{y}-h_{+}^{y}(h_{0}^{y})^{-1}h_{-}^{y}\end{pmatrix}$ (50)
$=$ $\det(h_{0}^{y})\det\begin{pmatrix}h_{0}^{y}&h_{+}^{y}&\cdots&0\\ h_{-}^{y}&h_{0}^{y}&\cdots&0\\ \vdots&\ddots&\ddots&\vdots\\ 0&\cdots&h_{-}^{y}&h_{0}^{y}\end{pmatrix}_{(L_{y}-1)\times(L_{y}-1)}$
@@
$H^{\prime}(k)=Q^{-1}H_{\text{y-OBC}}(k)Q,Q$ $=$ $\text{diag}\{1,r,r,r^{2},\cdots r^{L_{y}-1},r^{L_{y}-1},r^{L_{y}}\},$
$r$ $=$ $\sqrt{\left|\frac{(b_{0}-\cos k)^{B}}{a_{0}}\right|},$ (51)
@@
$H^{\prime}(k)=\begin{pmatrix}0&te^{-ik_{y}}+t^{\prime}\\ te^{ik_{y}}+t^{\prime}&0\end{pmatrix},$ (52)
where $t^{\prime}=\sqrt{a_{0}(b_{0}-\cos k)^{B}}$. Since a similarity matrix transform does not change the eigenspectrum, $H^{\prime}(k)$ possesses the same eigenvalues as $H_{\text{y-OBC}}$, with its bulk eigenvalues given by $E_{n}=\pm\sqrt{(te^{ik_{y}}+t^{\prime})(te^{-ik_{y}}+t^{\prime})}$. By substituting these values of $E_{n}$ into $\det[H_{\text{y-OBC}}(k)]=E_{e_{1}}(k)E_{e_{2}}(k)\prod_{n}E_{n}(k)$, we obtain
$\prod_{n}$ $E_{n}(k)$ $=-\prod_{k_{y}}(te^{ik_{y}}+t^{\prime})(te^{-ik_{y}}+t^{\prime})$ (53)
$\sim$ $t^{2L_{y}}+c_{1}t^{2L_{y}-1}t^{\prime}+c_{2}t^{2L_{y}-2}t^{\prime 2}+\cdots+c_{L_{y}}t^{\prime 2L_{y}},$
where in the RHS, the terms containing $e^{\pm ik_{y}}$ are incorporated into the coefficients $c_{1},c_{2},...,c_{L_{y}}$. As mentioned in the main text, the topological non-trivial condition is $|a_{0}(b_{0}-\cos k)^{B}|\leq t^{2}$, i.e. $|t^{\prime}|\leq|t|$. Consequently, the leading term in the RHS of the above equation is $t^{2L_{y}}$. Then using Eq (53), we have
$\det[H_{\text{y-OBC}}(k)]$ $=$ $E_{e_{1}}(k)E_{e_{2}}(k)*t^{2L_{y}}\left(1+O(t^{\prime}/t)\right)$
$=$ $\left(a_{0}(b_{0}-\cos k)^{B}\right)^{L_{y}}.$
$E_{e_{1}}(k)E_{e_{2}}(k)$$\approx$ $\frac{\left(a_{0}(b_{0}-\cos k)^{B}\right)^{L_{y}}}{t^{2L_{y}}}.$ (54)
@@
$\Delta=2|E_{e_{1}}|\sim\left(\frac{a_{0}(b_{0}-\cos k)^{B}}{t^{2}}\right)^{L_{y}/2}.$ (55)
II.2.2 Scaling of entanglement entropy $S_{A}$ with system sizes $L$ and $L_{y}$ in the case of gapless flat bands
In this subsection, we derive the scaling relations of $S_{A}$ as a function of $L_{y}$ and $L$, specifically Eq (13) in the main text:
$S_{A}\sim-\frac{1}{2}(BL_{y})^{2}\log L$ (56)
@@
Below, we present a comprehensive derivation of Eq (57). Under OBCs in the $y$ direction, $k$ is still a good eigenstate label and the projector operator in $k$ space is defined as $P(k)=\sum_{n\in occ}\ket{\psi_{n}(k)^{R}}\bra{\psi_{n}(k)^{L}}$ |. Due to the NHSE which pushes all right eigenstates in the y-direction towards the same edge, they become highly edge-localized and hence almost orthogonal. This large overlap ensures that the corresponding left edge eigenstates exhibit large amplitudes and contribute most significantly to $P(k)$. Therefore, $P(k)$ is dominated by the edge state contributions:
$P(k)$ $=$ $\sum_{n\in occ}\ket{\psi_{n}(k)^{R}}\bra{\psi_{n}(k)^{L}}$ | (58)
@@
In the following discussion, for the sake of brevity, we will omit the $k$ in |$\ket{\psi_{edge}(k)^{R}},\bra{\psi_{edge}(k)^{L}}$ |, rewriting them as |$\ket{\psi_{edge}^{R}},\bra{\psi_{edge}^{L}}$ |. To proceed, we note that $P_{edge}(k)$ is furthermore dominated by just one matrix element due to the exponential skin-localization of the edge states. To show this explicitly, we use the similarity transform [Eq (52)] to write the spatial profiles of the topological SSH edge states in the $y$ direction as
|$\ket{\psi^{\prime}_{edge}(y)}=\begin{pmatrix}\psi^{\prime A}(y)\\ \psi^{\prime B}(y)\end{pmatrix}\sim\begin{pmatrix}\epsilon^{y-1}\\ \epsilon^{L_{y}-y}\end{pmatrix},$
[…]
@@
$|\psi_{edge}^{R}\rangle=Q|\psi^{\prime}_{edge}\rangle\sim\begin{pmatrix}\begin{pmatrix}1\\ r\epsilon^{L_{y}-1}\end{pmatrix}\\ r\begin{pmatrix}\epsilon\\ r\epsilon^{L_{y}-2}\end{pmatrix}\\ \vdots\\ r^{L_{y}-2}\begin{pmatrix}\epsilon^{L_{y}-2}\\ r\epsilon\end{pmatrix}\\ r^{L_{y}-1}\begin{pmatrix}\epsilon^{L_{y}-1}\\ r\end{pmatrix}\end{pmatrix},$ (60)
and
$\bra{\psi_{edge}^{L}}=Q^{-1}\bra{\psi^{\prime}_{edge}}\sim\begin{pmatrix}\begin{pmatrix}1\\ \epsilon^{L_{y}-1}/r\end{pmatrix}\\ \frac{1}{r}\begin{pmatrix}\epsilon\\ \epsilon^{L_{y}-2}/r\end{pmatrix}\\ \vdots\\ \frac{1}{r^{L_{y}-2}}\begin{pmatrix}\epsilon^{L_{y}-2}\\ \epsilon/r\end{pmatrix}\\ \frac{1}{r^{L_{y}-1}}\begin{pmatrix}\epsilon^{L_{y}-1}\\ 1/r\end{pmatrix}\end{pmatrix}^{T}.$ (61)
@@
$P(k)\approx P_{\text{edge}}(k)=\ket{\psi_{\text{edge}}^{R}}\bra{\psi_{\text{edge}}^{L}}$ | (62)
$\sim$
$\begin{pmatrix}\begin{pmatrix}1&\epsilon^{L_{y}-1}/r\\ r\epsilon^{L_{y}-1}&\epsilon^{2L_{y}-2}\end{pmatrix}&\frac{1}{r}\begin{pmatrix}\epsilon&\epsilon^{L_{y}-2}/r\\ r\epsilon^{L_{y}}&\epsilon^{2L_{y}-3}\end{pmatrix}&\cdots&\color[rgb]{0,0,1}\frac{1}{r^{L_{y}-1}}\begin{pmatrix}\epsilon^{L_{y}-1}&1/r\\ r\epsilon^{2L_{y}-2}&\epsilon^{L_{y}-1}\end{pmatrix}\\ r\begin{pmatrix}\epsilon&\epsilon^{L_{y}}/r\\ r\epsilon^{L_{y}-2}&\epsilon^{L_{y}-2}\end{pmatrix}&\begin{pmatrix}\epsilon^{2}&\epsilon^{L_{y}-1}/r\\ r\epsilon^{L_{y}-1}&\epsilon^{2L_{y}-4}\end{pmatrix}&\cdots&\frac{1}{r^{L_{y}-2}}\begin{pmatrix}\epsilon^{L_{y}}&\epsilon/r\\ r\epsilon^{2L_{y}-3}&\epsilon^{L_{y}-2}\end{pmatrix}\\ \vdots&\vdots&\vdots&\vdots\\ r^{L_{y}-1}\begin{pmatrix}\epsilon^{L_{y}-1}&\epsilon^{2L_{y}-2}/r\\ r&\epsilon^{L_{y}-1}\end{pmatrix}&r^{L_{y}-2}\begin{pmatrix}\epsilon^{L_{y}}&\epsilon^{2L_{y}-3}/r\\ r\epsilon&\epsilon^{L_{y}-2}\end{pmatrix}&\cdots&\begin{pmatrix}\epsilon^{2L_{y}-2}&\epsilon^{L_{y}-1}/r\\ r\epsilon^{L_{y}-1}&1\end{pmatrix}\end{pmatrix}_{2L_{y}\times 2L_{y}}$
@@
The truncated projector $\bar{P}$ can be obtained by Fourier transforming each matrix element on the right-hand side (RHS) of Eq (62), i.e.,
$\bar{P}=\begin{pmatrix}\bar{P}_{1,1}&\bar{P}_{1,2}&\cdots&\bar{P}_{1,2L_{y}}\\ \bar{P}_{2,1}&\bar{P}_{2,2}&\cdots&\bar{P}_{2,2L_{y}}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{P}_{2L_{y},1}&\bar{P}_{2L_{y},2}&\cdots&\bar{P}_{2L_{y},2L_{y}}\\ \end{pmatrix},$ (63)
where $\bar{P}_{m,n}$ is an $l\times l$ matrix whose elements are
$\bra{x_{1}}\bar{P}_{m,n}\ket{x_{2}}=\frac{1}{L}\sum_{k}e^{ik(x_{1}-x_{2})}P_{m,n}(k),$ (64)
@@
with $b_{0}=1$ and the leading $k=\pm k_{1}=\pm\pi/L$ contributions substituted to obtain the 3rd line. The other matrix elements scale more slowly with $L$. Since $\bra{x_{1}}\bar{P}_{2L_{y},1}\ket{x_{2}}\sim O(1)$, we have established Eq (57):
$\text{Tr}(\bar{P}^{2})\approx 2\text{Tr}(\bar{P}_{1,2L_{y}}\bar{P}_{2L_{y},1})\sim c_{1}L^{BL_{y}-1}+c_{2}L^{BL_{y}-1-2}+\cdots$ (66)
@@
$S_{A}(p_{i})$ $=$ $-p_{i}\log p_{i}-(1-p_{i})\log(1-p_{i})$ (67)
$=$ $-cL^{\alpha}\log(cL^{\alpha})-(1-cL^{\alpha})\log(1-cL^{\alpha})$
@@
$=$ $(cL^{\alpha}+1)\log(-1)-\log(cL^{\alpha})$
$\approx$ $-{\color[rgb]{0,0,1}\alpha\log L}+cL^{\alpha}\pi i-\log c.$
The other eigenvalue with $p_{i}^{\prime}=1-p_{i}$ (see subsection C.2) contributes the same real part to $S_{A}$, but opposite imaginary part that cancels off, as numerically observed as twofold degenerate states. Therefore, the total entanglement entropy is dominated by the $\log L$ contribution in blue above which should be multiplied by 4, i.e.,
$S_{A}$ $=$ $\sum_{p_{i}}-{p_{i}}\log p_{i}-(1-p_{i})\log(1-p_{i})$ (68)
$\sim$
$-4\times\left[\frac{BL_{y}-1}{2}+\left(\frac{BL_{y}-1}{2}-1\right)+\left(\frac{BL_{y}-1}{2}-2\right)+\cdots\right]\log L.$
@@
$=$ $-\left\lceil\frac{B^{2}L_{y}^{2}-1}{2}\right\rceil\log L\quad\sim-\frac{1}{2}(BL_{y})^{2}\log L,$
@@
Figure 7: Excellent agreement of our super-volume-law negative entanglement scaling result with numerics for various $B$ and $L_{y}$, both for the $\bar{P}$ eigenvalues $p_{i}$ [Eq (57)] (a-c) and the corresponding entanglement scaling $S_{A}\sim-\left\lceil(B^{2}L_{y}^{2}-1)/2\right\rceil\log L$ [Eq (69)]. Parameters: $t=0.5,a_{0}=2,b_{0}=1$.
Below, we further present the numerical verification of Eq (68). In Fig. 7(a), we take $L_{y}=3,B=1$, since $\frac{BL_{y}-1}{2}=1$, only the first order $p_{i}\sim L^{1}$ exists. The numerical fitting shown in Fig. 7(d) yields:
$S_{A}\sim-4.4072\log L\quad\approx-4\times\frac{BL_{y}-1}{2}\log L=-4\log L,$ (70)
which demonstrates good agreement with Eq (68). In Fig. 7(b), where $L_{y}=5,B=1$, $\frac{BL_{y}-1}{2}=2$, both the first and second orders of $L$ in $p_{i}\sim L^{1}$, $p_{i}\sim L^{2}$ are present. We numerically obtain the scaling of entanglement entropy as depicted in Fig. 7(e):
$S_{A}$ $\sim-11.7648\log L$ (71)
$\approx$ $-4\times\left(\frac{BL_{y}-1}{2}+\frac{BL_{y}-1}{2}-1\right)\log L=-12\log L.$
For the case of $L_{y}=3,B=2$ shown in Fig. 7(c), the first, second and third orders of $L$ all contribute to the EE:
$S_{A}$ $\sim$ $-17.1980\log L$ (72)
$\approx$ $-4\times\left(\frac{BL_{y}-1}{2}+\frac{BL_{y}-1}{2}-1+\frac{BL_{y}-1}{2}-2\right)\log L$
$=$ $-18\log L.$
@@
Figure 8: Negative entanglement scaling for the gapped EP case ($b_{0}>1$). (a) As numerically plotted for $L=9,b_{0}=1.2$, the approximate linear scaling of the scaling exponent $M$ of the dominant $\bar{P}$ eigenvalue with $L_{y}$, as described by Eq (75). (b) The negatively linear relationship of the entanglement entropy $S_{A}$ with $\log L$ for different $L_{y}$ ranging from 22 to 32, with $L\in[22,32]$ and parameters $t=0.8,a_{0}=1,b_{0}=1.2,B=1$. As predicted in Eq (78), the gradient indeed increases with $L_{y}$. (c) Using the same parameters as in (b), approximate linear dependence of the gradient of $S_{A}$ on $L_{y}$. (d) For large $L=60$ and parameters $t=0.8,B=1,b_{0}=1.2,a_{0}=2$, $S_{min}$ indeed scales linearly with $L_{y}$, as predicted in Eq (81).
II.2.3 Scaling of entanglement entropy $S_{A}$ for gapped flat bands
In the case of gapped bands where $b_{0}>1$ (e.g., $b_{0}=1.2$), there exists a small gap $\Delta$ in the topological flat edge bands. Due to this gap, the entanglement entropy $S_{A}$ decreases with increasing system size $L$ and eventually stabilizes at a minimum constant value, $S_{min}$. In this section, we will demonstrate that: